Simple math problem (which I can not solve). Word for word with no additional information:
Red beads are worth 1 cent. Blue beads are worth 10 cents. Yellow beads are worth 100 cents. I have 1 yellow bead, 2 blue beads, and 3 red beads. I owe Amani 4 blue and 5 red beads. How can I pay him?

Answer by JBarnum(2146)

R=1 B=10 Y=100
you have Y+2B+3R=123cents
you owe 4B+5R=45cents
so your gonna have to trade that yellow bead in for 10 blue beads cause 10B=100cents
so now u have 12B+3R=123cents, now trade 1 Blue for 10Red beads since 1B=10 cents
so now you have 11B+13R=123cents
now give Amani 4Blue beads and 5Red beads leaving u with 7B+8R=78cents
Check by 123-45=78. correct